The given information is available for two samples selected from independent normally distributed populations. Population A: n 1=24​, Upper S Subscript 1 Superscript 2=166.6 Population B: n 2=24​, Upper S Subscript 2 Superscript 2=136.4 How many degrees of freedom are there in the numerator and denominator of the F​ test? There are ____ degrees of freedom in the numerator. There are ___degrees of freedom in the denominator.
Answer #1

Ans. we have n,=24, n2 = 24 Numerator degrees of freedom = 1,-1=24-1 = 23 Denominator degrees of freedom = n2 = 24-1